NOTES

(a) If you want to get TeX files instead of LaTeX files, remove ‘%’ from the first line of file ‘Exam-

ples.tex’.

(b) If you want different extensions in the files’ names, change the definition of \ext in the second line

of file ‘Examples.tex’.

DraTeX

\underline{tex sec62a.Changeable} First example in section 6.2 (for TeX)

\input DraTex.sty
\Draw
\LineAt(0.1,  0,-18,-10)
\LineAt(-18,-10,- 6, 26)
\LineAt(- 6, 26, 18,- 4)
\LineAt( 18,- 4, 12,-40)
\LineAt( 18,- 4, 24,-40)
\LineAt( 18,- 4, 58,- 4)
\LineAt( 58,- 4, 52,-40)
\LineAt(58,-4,64,-40) \LineAt(58,-4,66,4)
\EndDraw
\bye

\underline{latex sec62b.Changeable} First example in section 6.2 (for LaTeX)

\documentstyle{report}
\input DraTex.sty
\begin{document}
\Draw
\LineAt(0.1,  0,-18,-10)
\LineAt(-18,-10,- 6, 26)
\LineAt(- 6, 26, 18,- 4)
\LineAt( 18,- 4, 12,-40)
\LineAt( 18,- 4, 24,-40)
\LineAt( 18,- 4, 58,- 4)
\LineAt( 58,- 4, 52,-40)
\LineAt(58,-4,64,-40) \LineAt( 58,-4,66,4)
\EndDraw
\end{document}

\underline{latex sec62c.Changeable} First example in section 6.2—second version (for LaTeX)

\documentstyle[DraTex]{report}
\begin{document}
\Draw
\LineAt(0.1,  0,-18,-10)
\LineAt(-18,-10,- 6, 26)
\LineAt(- 6, 26, 18,- 4)
\LineAt( 18,- 4, 12,-40)
\LineAt( 18,- 4, 24,-40)
\LineAt( 18,- 4, 58,- 4)
\LineAt( 58,- 4, 52,-40)
\LineAt(58,-4,64,-40) \LineAt(58,-4,66,4)
\EndDraw
\end{document}

\underline{latex exr672.Changeable} Exercise 6.7.2 (for LaTeX)

 \MarkLoc(a)    \Move(50,0)
 \MarkLoc(A)    \Move(50,0)
 \MarkLoc(b)    \Move(-30,30)
 \MarkLoc(B)    \Move(-30,30)
 \MarkLoc(c)    \Move(-20,-30)
 \MarkLoc(C)
 \Curve(A,b,b,B)
 \Curve(B,c,c,C)
 \Curve(C,a,a,A)

\underline{tex sec61.Changeable} First example in section 6.1 (for TeX)

\input DraTex.sty
We can draw clocks of
different styles and sizes.
\Draw
\Move(0,20)     \Line(0,-50) \Line(50,0)
\PenSize(1.5pt) \Line(0,50)  \Line(-50,0)
\Move(5,-25)    \Text(--9--)
\Move(20,-18)   \Text(--6--)
\Move(20,18)    \Text(--3--)
\Move(-20,18)   \Text(--12--)
\Move(0,-18)    {\RotateTo(60) \LineF(19)}
\RotateTo(210)  \LineF(14)
\EndDraw
We can draw people of
different shapes and looks.
\Draw(0.6pt,0.6pt)
\DrawOvalArc(30,45)(0,180)
\DrawOvalArc(30,12)(180,360)
\DrawOvalArc(60,25)(140,400)
\Move(0,5)
\DrawOvalArc(30,60)(225,315)
\DrawOvalArc(20,52)(260,280)
\Move(-10,-37) {\Text(--$\cdotp$--)}  \Text(--o--)
\Move(20,0)    {\Text(--$\cdotp$--)}  \Text(--o--)
\EndDraw
And we can produce many other
types of drawings.       \bye

\underline{latex sec61a.Changeable} First example in section 6.1 (for LaTeX)

\documentstyle{report}
\input DraTex.sty
\begin{document}

We can draw clocks of
different styles and sizes.
\Draw
\Move(0,20)     \Line(0,-50) \Line(50,0)
\PenSize(1.5pt) \Line(0,50)  \Line(-50,0)
\Move(5,-25)    \Text(--9--)
\Move(20,-18)   \Text(--6--)
\Move(20,18)    \Text(--3--)
\Move(-20,18)   \Text(--12--)
\Move(0,-18)    {\RotateTo(60) \LineF(19)}
\RotateTo(210)  \LineF(14)
\EndDraw
We can draw people of
different shapes and looks.
\Draw(0.6pt,0.6pt)
\DrawOvalArc(30,45)(0,180)
\DrawOvalArc(30,12)(180,360)
\DrawOvalArc(60,25)(140,400)
\Move(0,5)
\DrawOvalArc(30,60)(225,315)
\DrawOvalArc(20,52)(260,280)
\Move(-10,-37) {\Text(--$\cdotp$--)}  \Text(--o--)
\Move(20,0)    {\Text(--$\cdotp$--)}  \Text(--o--)
\EndDraw
And we can produce many other
types of drawings.       \end{document}

\underline{latex sec61b.Changeable} First example in section 6.1—second version (for LaTeX)

\documentstyle[DraTex]{report}
\begin{document}
We can draw clocks of
different styles and sizes.
\Draw
\Move(0,20)     \Line(0,-50) \Line(50,0)
\PenSize(1.5pt) \Line(0,50)  \Line(-50,0)
\Move(5,-25)    \Text(--9--)
\Move(20,-18)   \Text(--6--)
\Move(20,18)    \Text(--3--)
\Move(-20,18)   \Text(--12--)
\Move(0,-18)    {\RotateTo(60) \LineF(19)}
\RotateTo(210)  \LineF(14)
\EndDraw
We can draw people of
different shapes and looks.

\Draw(0.6pt,0.6pt)
  \DrawOvalArc(30,45)(0,180)
  \DrawOvalArc(30,12)(180,360)
  \DrawOvalArc(60,25)(140,400)
  \Move(0,5)
  \DrawOvalArc(30,60)(225,315)
  \DrawOvalArc(20,52)(260,280)
  \Move(-10,-37) {\Text(--$\cdotp$--)}  \Text(--o--)
  \Move(20,0)    {\Text(--$\cdotp$--)}  \Text(--o--)
\EndDraw
And we can produce many other
types of drawings.       \end{document}

\underline{latex exr732.Changeable} Exercise 7.3.2 (for LaTeX)

\font\ARROWS=line10 scaled\magstep5
\font\CIRCLES=lcircle10 scaled\magstep5
%   \newfont{\ARROWS}{line10 scaled\magstep5}
%   \newfont{\CIRCLES}{lcircle10 scaled\magstep5}
                             \Text(--\CIRCLES \char 10--)
\Text(--\CIRCLES \char 11--) \Text(--\ARROWS \char 55--)
\Text(--\CIRCLES \char 110--)\EntryExit(-1,0,-1,0)
\Text(--\CIRCLES \char 114--)

\underline{latex exr741.Changeable} Exercise 7.4.1 (for LaTeX)

\Ragged(8)
\Line(10,-10) \Line(10,20)
\LineTo(0,0)

\underline{latex sec81.Changeable} First example in section 8.1 (for LaTeX)

\ThreeDim(-150,-60,-600)
   \Line(60,0,0)   \Line(0,0,60)
   \Line(-60,0,0)  \Line(0,0,-60)
   \LineTo(0,60,0) \Line(60,0,0)
   \Line(0,-60,0)
   \LineAt(0,60,0,0,60,60)
   \Line(0,-60,0)
\EndThreeDim

\underline{latex usa.Changeable} Example of USA map in section 9.3 (for LaTeX)

 (1.4pt,1.4pt)
\Table\USA{       % California
   -60,-35 & -77,-30 & -80,
   -20 & -90,-15 & -100,40 &
   % Oregon
   -100,50 & -90,80 &

% Wash
-90,95 & -78,90 & -75,95 &
% Minn
100,70 & 85,60 &
% Wisc-Mich
95,60 & 105,65 & 102,63 & 110,60 & 125,60 & 110,55 & 110,30 &
115,25 & 120,30 & 115,40 & 115,50 & 120,55 & 130,55 & 140,40 &
135,25 &
% Ohio-NY-Vt
145,25 & 155,30 & 160,45 & 175,45 & 175,55 & 180,60 & 195,65 &
% Maine
200,75 & 200,85 & 205,90 & 210,90 & 215,80 & 220,75 & 205,60 & 205,50 &
% Mass--NY
210,43 & 208,47 & 215,43 & 190,30 & 202,35 &
% NJ--Va
197,25 & 190,18 & 185,20 & 190,15 & 190,5 & 182,18 & 182,0 & 185,0 &
% NC
190,-8 & 185,-4 & 183,-6 & 186,-6 & 182,-15 &
190,-18 & 190,-10 & 192,-10 &
195,-18 &
% Ga-Fl
165,-40 & 160,-50 & 175,-80 & 175,-90 & 170,-100 & 165,-100 &
155,-80 & 155,-65 & 145,-60 & 140,-62 & 136,-62 & 120,-60 &
% Alaba--La
102,-63 & 107,-67 & 107,-72 & 110,-75 & 90,-65 &
78,-65 &
% Texas
62,-78 & 57,-78 & 50,-82 & 53,-92 & 48,-90 & 25,-65 & 15,-65 &
15,-68 & -10,-50 &
% New Mexico
-40,-50 &
% Back to California
-60,-35 }
{ \USA(0,0){\MoveTo} \USA(1,99){\LineTo} } \Text(--USA--)
\Scale(0.5,0.5) \USA(46,46){\MoveTo} \USA(47,92){\LineTo}
\MoveTo(-30,0) \MarkLoc(a) \USA(0,0){\Move} \MarkLoc(b)
\USA(0,46){{\LineToLoc(b)} \MarkLoc(b) \MoveToLoc(a) \Move}

\underline{latex exr1011a.Changeable} Exercise 10.1.1 (for LaTeX)

\Text(--Government~%
          Bonds (\%)--)
\Table\B{Britain,9.59 &
Canada,7.98 &  Japan,5.46 &
U.S.,  6.84 &  Mexico,14.9}
\B(0,4){\PutBar}

\underline{latex exr1011b.Changeable} Exercise 10.1.1 (for LaTeX)

\chick(0,75)  \Scale(-1,1)

\chick(0,140)

\underline{latex exr1011c.Changeable} Exercise 10.1.1 (for LaTeX)

\Table\x{ 30,40
  & 15,60 & 27,15 & 35,35
  & 50,45 & 25,70 & 20,20
  & 25,75 }   \x(0,99){\Tr}

\underline{latex exr1011d.Changeable} Exercise 10.1.1 (for LaTeX)

\sqr(-24,0)
\sqr(0,32)
\sqr(24,-32)

\underline{latex exr1011e.Changeable} Exercise 10.1.1 (for LaTeX)

\engine(90)  \MoveTo(50,0)
\engine(180) \MoveTo(100,0)
\engine(-30)

\underline{latex exr1011f.Changeable} Exercise 10.1.1 (for LaTeX)

\MarkLoc(A)
\spring(110) \Rotate(70)
\spring(60)  \MarkLoc(B)
\DSeg\RotateTo(B,A)
\LSeg\spring(B,A)

\underline{latex exr1021.Changeable} Exercise 10.2.1 (for LaTeX)

\new(BODY) \<BODY>(20)
\LineF(20) \Rotate(55) \LineF(16)
\new(hand) \<BODY>(140) \LineF(35)
\new(hand) \<BODY>(90) \LineF(7)
\new(head) \<BODY>(-60) \LineF(19)
\new(left-leg) \<BODY>(210) \LineF(12)
\Rotate(30) \LineF(14) \new(right-leg)
\Do(1,11){\I+15; \<head>(\Val\I)
\LineF(11)}

\underline{latex exr1022.Changeable} Exercise 10.2.2 (for LaTeX)

\def\DefineSymbol#1{%
  \Indirect\Define<#1>}
\def\DrawText(--#1--){%
    \let\xNextSym=\xDrawSym  \xNextSym#1{}}
\def\xDrawSym#1{\def\temp{#1}%
    \ifx \temp\empty  \let \xNextSym=\relax
    \else \Indirect<#1>\fi   \xNextSym }
\DefineSymbol A{{ \Line(21,60)
  \Line(21,-60)  \Move(-7,20)
  \Line(-28,0) } \Move(52,0)  }
\DefineSymbol B{{ \DrawText(--P--)
  \Move(-25,18)
  \DrawOvalArc(15,18)(-90,90)
  \Move(0,-18)   \Line(-25,0)}
  \Move(50,0) }    \DefineSymbol P
{{      \Line(0,60)    \Line(25,0)
    \Move(0,-12) \DrawOvalArc(15,12)(-90,90)
    \Move(0,-12)  \Line(-25,0) }      \Move(50,0)  }
\DefineSymbol C{{     \Move(20,30)
    \DrawOvalArc(20,30)(45,315) }    \Move(46.6,0)  }
\DefineSymbol D{{    \Line(0,60)    \Line(20,0)
    \Move(0 ,-30) \DrawOvalArc(20,30)(-90,90) \Move(0,-30)
    \Line(-20,0) }   \Move(50,0)  }
\DefineSymbol E{{  \DrawText(--F--)  \Move(-50,0)
    \Line(40,0) }  \Move(50,0)      }
\DefineSymbol F{{   \Line(0,60)   \Line(40,0)
    \Move(-40,-30) \Line(30,0)  }  \Move(50,0)     }
\DefineSymbol G{{ \DrawText(--C--)
   \Move(-10,30) {\Line(-20,0)} \Line(0,-30)   }
   \Move(46.6,0)     }
\DefineSymbol H{{ \Line(0,60)  \Move(40,0)
   \Line(0,-60)  \Move(0,30)  \Line(-40,0)   }  \Move(50,0)     }
\DefineSymbol I{{ \Line(0,60)  }  \Move(10,0)     }
\DefineSymbol J{{ \Move(20,17) \DrawOvalArc(20,17)(180,360)
   \Move(20,0) \Line(0,43)   }  \Move(50,0)     }
\DefineSymbol K{{ \Line(0,60)  \Move(0,-35)
   \Line(40,35) \Move(-32,-28) \Line(32,-32) }  \Move(50,0)     }
\DefineSymbol L{{ { \Line(0,60) } \Line(40,0)  }  \Move(50,0)     }
\DefineSymbol M{{ \Line(0,60)  \Line(25,-60)
   \Line(25,60) \Line(0,-60)   }  \Move(60,0)     }
\DefineSymbol N{{ \Line(0,60) \Line(40,-60)
   \Line(0,60)  }  \Move(50,0)     }
\DefineSymbol O{{ \Move(22,30)  \DrawOval(22,30)   }  \Move(54,0)     }
\DefineSymbol P{{      \Line(0,60)    \Line(25,0)
    \Move(0,-12) \DrawOvalArc(15,12)(-90,90)
    \Move(0,-12)  \Line(-25,0) }      \Move(50,0)  }
\DefineSymbol Q{{ \DrawText(--O--)  \Move(-10,-5)
    \Line(-15,15)   }  \Move(55,0)     }
\DefineSymbol R{{ \DrawText(--P--) \Move(-10,0)
    \Line(-15,36)   }  \Move(50,0)     }

\DefineSymbol S{{  \Move(20,18) \DrawOvalArc(20,18)(-140,90)
 \Move(0,30)  \DrawOvalArc(20,12)(40,270)  }  \Move(50,0)     }
\DefineSymbol T{{ \Move(22,0) \Line(0,60)  \Move(-22,0)
 \Line(44,0)   }  \Move(54,0)     }
\DefineSymbol U{{ \Move(20,15)  \DrawOvalArc(20,15)(180,360)
 \Move(-20,0) \Line(0,45) \Move(40,0) \Line(0,-45)   }
 \Move(50,0)     }
\DefineSymbol V{{ \Move(0,60) \Line(22,-60)
 \Line(22,60)   }  \Move(54,0)     }
\DefineSymbol W{{  \Move(0,60) \Line(15,-60)
 \Line(15,60)  \Line(15,-60) \Line(15,60)  }  \Move(70,0)     }
\DefineSymbol X{{ \Line(40,60)  \Move(-40,0)
 \Line(40,-60)   }  \Move(50,0)     }
\DefineSymbol y{{ \Move(20,0) \Line(0,34) { \Line(-20,26)}
 \Line(20,26)   }  \Move(50,0)     }
\DefineSymbol Z{{ \Move(0,60)  \Line(40,0) \Line(-40,-60)
 \Line(40,0)   }  \Move(50,0)     }
\DefineSymbol 0{{ \Move(25,30) \DrawOval(25,30)   }  \Move(60,0)     }
\DefineSymbol 1{{ \Line(0,60)   }  \Move(10,0)     }
\DefineSymbol 2{{ \MarkLoc(a) \Move(0,10)
\MarkLoc(a')  \Move(40,30)  \MarkLoc(b)
\Move(0,-10)  \MarkLoc(b')  \Curve(a,a',b',b)
\Move(-20,10) \DrawOvalArc(20,20)(0,180)
\Move(-20,-40) \Line(40,0)   }  \Move(50,0)     }
\DefineSymbol 3{{ \MarkLoc(a')  \Move(0,10)  \MarkLoc(a)
\Move(20,-10) \MarkLoc(b) \Move(-10,0) \MarkLoc(b')
\Curve(a,a',b',b)  \Move(22,0) \MarkLoc(b')
\Move(8,8)  \MarkLoc(a')  \Move(0,10) \MarkLoc(a)
\Curve(a,a',b',b)  \Move(0,10)  \MarkLoc(a')
\Move(-8,8) \MarkLoc(b') \Move(-10,0)
\MarkLoc(b) \Curve(a,a',b',b) \Move(-5,0)
\MarkLoc(b') \Move(-5,-3) \MarkLoc(a')
\Move(0,-2)  \MarkLoc(a)  \Curve(a,a',b',b)
\Line(30,29) \Line(-40,0)   }  \Move(50,0)     }
\DefineSymbol 4{{      \Move(35,0)  \Line(0,60)
 \Line(-35,-35)     \Line(45,0)}    \Move(55,0)  }
\DefineSymbol 5{{ \Move(15,20) \DrawOvalArc(19,20)(-140,140)
\Move(-15,10) \Line(0,30) \Line(30,0)   }  \Move(44,0)     }
\DefineSymbol 6{{ \Move(22,30)  \DrawOvalArc(22,30)(70,270)
\Move(0,-10) \DrawOvalArc(22,15)(0,174)
\DrawOvalArc(22,20)(270,360)  }  \Move(54,0)     }
\DefineSymbol 7{{ \Move(0,60) \Line(40,0)
\MarkLoc(a) \Move(-20,-60) \MarkLoc(b)
\DSeg\RotateTo(b,a) \Rotate(20) \MoveF(10) \MarkLoc(b')
\MoveToLoc(a)
\DSeg\RotateTo(a,b) \Rotate(-20) \MoveF(10) \MarkLoc(a')
\Curve(a,a',b',b)  }  \Move(50,0)     }
\DefineSymbol 8{{  \Move(20,17) \DrawOval(20,17)
\Move(0,30)  \DrawOval(20,13)  }  \Move(50,0)     }
\DefineSymbol 9{{\Move(50,60) \Scale(-1,-1) \DrawText(--6--)   }  \Move(54,0)
}
\Scale(0.1,0.1) \Ragged(1)

\DrawText(--BOLD--)
\PenSize(0.2pt) \MoveTo(0,-100)
{ \Scale(-1,1)
\DrawText(--BACKWARD--) }
\MoveTo(0,-200) \RotatedAxes(0,75)  \DrawText(--SLANTED--)
\EndRotatedAxes \MoveTo(0,-300) \RotatedAxes(-45,45)
\DrawText(--ROTATED--)          \EndRotatedAxes

\underline{latex boxes.Changeable} Last example in section 11.2 (for LaTeX)

\Object\ch(1){\Text(--#1--)
\MoveToExit(-1,-1)
\FigSize\Q\R{\Text(--#1--)}
\DrawRect(\Val\Q,\Val\R) }
\Define\chrs(1){\font\f=#1\space scaled\magstep4 \f
\Table\chrs{A & l & l & ~ & i & n & ~ &
    b & o & x & e & s & .}   \chrs(0,99){\ch} }
\PenSize(0.2pt) \EntryExit(-1,-1,1,-1)
{\chrs(cmtt10)} \Move(0,-20) \chrs(cmsl10)

\underline{latex exr1121.Changeable} Exercise 11.2.1 (for LaTeX)

\Define\side(3){
\ProjectedAxes(#1,#2)(#1,#3)
  \DrawRect(1,1)  \Move(0.5,0.5)
  \DrawCircle(0.5)
\EndProjectedAxes }
\ThreeDim(80,70,-40)    \MarkPLoc(a)
\Move(0,-50,0)   \MarkPLoc(b)
\Move(0,50,25)   \MarkPLoc(c)
\Move(-25,0,-25) \MarkPLoc(d)
\side(a,b,d) \side(a,b,c) \side(a,c,d)
\EndThreeDim

\underline{latex exr1124.Changeable} Exercise 11.2.4 (for LaTeX)

\Define\Spline(1){
\MarkLoc(0) \MoveTo(0,0) \MarkLoc(00)
\Table\P{#1} \Define\Count(1){\K+1;} \K=-4;
\P(0,999){\Count}
\Do(0,\Val\K){
  \GetCoefficients
  \MoveTo(\Val\Ao,\Val\Bo)  \CSeg\Move(00,0)
  \MarkLoc(o)   \f   \MarkLoc(x)
  \Do(1,30){
     \T=\DoReg;  \T/30;
     \X=\Aiii; \X*\T; \X+\Aii; \X*\T; \X+\Ai; \X*\T;
     \Y=\Biii; \Y*\T; \Y+\Bii; \Y*\T; \Y+\Bi; \Y*\T;

    \MoveToLoc(o)    \Move(\Val\X,\Val\Y)      \f
    {\LineToLoc(x)}  \MarkLoc(x)    }}
\MoveToLoc(0)}
\Define\f{}
\DecVar\Ao \DecVar\Ai \DecVar\Aii \DecVar\Aiii \DecVar\X
\DecVar\Bo \DecVar\Bi \DecVar\Bii \DecVar\Biii \DecVar\Y
\Define\GetCoefficients{
\I=\DoReg;  \P(\Val\I,\Val\I){\First}
\I+1;       \P(\Val\I,\Val\I){\Second}
\I+1;       \P(\Val\I,\Val\I){\Third}
\I+1;       \P(\Val\I,\Val\I){\Fourth}  }
\Define\First(2){\Ao=#1; \Ai=-#1; \Aii=#1; \Aiii=-#1;
              \Bo=#2; \Bi=-#2; \Bii=#2; \Biii=-#2;}
\Define\Second(2){
\T=#1; \T* 4; \Ao +\T;   \T=#2; \T* 4; \Bo +\T;
\T=#1; \T*-2; \Aii+\T;   \T=#2; \T*-2; \Bii+\T;
\T=#1; \T* 3; \Aiii +\T; \T=#2; \T* 3; \Biii +\T;}
\Define\Third(2){ \Ao +#1; \Ao /6;   \Bo +#2; \Bo /6;
               \Ai +#1; \Ai /2;   \Bi +#2; \Bi /2;
               \Aii+#1; \Aii/2;   \Bii+#2; \Bii/2;
\T=#1; \T*-3; \Aiii+\T;  \T=#2; \T*-3; \Biii+\T;}
\Define\Fourth(2){
\Aiii+#1;  \Aiii/6;  \Biii+#2; \Biii/6;}
\Spline(0,15 & 30,-70
& 40,-50 & 10,0 &  15,30
& 50,5 &60,10 & 8,45 &
8,55 & 30,70 &  0,80 &
-30,70 &-8,55 & -8,45 &
-60,10 &-50,5 & -15,30 &
-10,0 &-40,-50 &-30,-70 &
0,15 & 30,-70 & 40,-50 )
\Define\ShowPoint(2){
\MoveToLoc(0) \Move(#1,#2)
\Text(--$\circ$--)}
\P(0,99){\ShowPoint}
\Define\TEX{
\Spline(0,-15 & 0,-15 &
       0,-15 & 0, 15 &
       0, 15 & 0, 15 )
\Spline(-10,15 &
  -10,15 & -10,15 & 10,15 &
   10,15 &  10,15)
\Move(12,0)
0,0 &  0,0 & 14,0 & 14,0 & 14,0)
\Spline(  20,15  &20,15  &20,15  & 0,15 & 0,15
&   0,-15 & 0,-15 & 20,-15 &20,-15&20,-15)
\Move(32,0) \Spline(  -10,-15  &  -10,-15  &
-10,-15 & 10,15 & 10,15 & 10,15)
\Spline(  -10,15  &

   -10,15 &    -10,15 & 10,-15 & 10,-15 & 10,-15) }
\TEX \MoveTo(0,0)   \F
\TEX \MoveTo(0,-35) \G
\TEX

\underline{latex exr1131.Changeable} Exercise 11.3.1 (for LaTeX)

\Define\put(2){
   \IF \EqText(,#1) \THEN
      \top  \Define\put(1){}
   \ELSE \push(#1) \FI  \put(#2)}
\Define\stack(1){
   {\put(#1,,)} \Move(0,-13)}
\Table\x{ &  1&  1, 2&
1, 2, 3&        1, 2, 3, 4&
1, 2, 3, 4, *&  1, 2, 12&
1, 2, 12,+&  1, 14&  1, 14,+& 15}
\x(0,100){\stack}

\underline{latex exr1132.Changeable} Exercise 11.3.2 (for LaTeX)

\Table\g{
  16 & 8 & 4 & 2 & 1 &
  20 & 16 & 8 & 4 & 2 &
  17 & 12 & 16 & 8 & 4 &
  23 & 20 & 12 & 13 & 17 }
\ThreeDim(160,75,-70)
  \Scale(20,5,20)
  \g(0,140){\surface}
\EndThreeDim

AlDraTeX

\underline{tex sec132.Changeable} Example in section 13.2 (for TeX)

\input DraTex.sty
\input AlDraTex.sty
\Draw
\Scale(1,0.5)
\PieChartSpec(1,50,20)()
\PieChart(15 & 10 & 30)
\Move(0,-6)
\DrawOvalArc(50,50)(180,360)
\EndDraw
\bye

\underline{tex sec132a.Changeable} Example in section 13.2—second version (for TeX)

\input DraTex.sty
\def\AlDraTex{pie}
\input AlDraTex.sty
\Draw
\Scale(1,0.5)
\PieChartSpec(1,50,20)()
\PieChart(15 & 10 & 30)
\Move(0,-6)
\DrawOvalArc(50,50)(180,360)
\EndDraw
\bye

\underline{latex sec132b.Changeable} Example in section 13.2 (for LaTeX)

\documentstyle{report}
\input DraTex.sty
\input AlDraTex.sty
\begin{document}
\Draw
\Scale(1,0.5)
\PieChartSpec(1,50,20)()
\PieChart(15 & 10 & 30)
\Move(0,-6)
\DrawOvalArc(50,50)(180,360)
\EndDraw
\end{document}

\underline{latex sec132c.Changeable} Example in section 13.2—second version (for LaTeX)

\documentstyle[DraTex,AlDraTex]{report}
\begin{document}
\Draw
\Scale(1,0.5)
\PieChartSpec(1,50,20)()
\PieChart(15 & 10 & 30)
\Move(0,-6)
\DrawOvalArc(50,50)(180,360)

\EndDraw
\end{document}

\underline{latex sec132d.Changeable} Example in section 13.2—third version (for LaTeX)

\documentstyle{report}
   \input DraTex.sty
      \def\AlDraTex{pie}
   \input AlDraTex.sty
\begin{document}
\Draw
   \Scale(1,0.5)
   \PieChartSpec(1,50,20)()
   \PieChart(15 & 10 & 30)
   \Move(0,-6)
   \DrawOvalArc(50,50)(180,360)
\EndDraw
\end{document}

\underline{latex sec132e.Changeable} Example in section 13.2—forth version (for LaTeX)

\def\AlDraTex{pie}
\documentstyle[DraTex,AlDraTex]{report}
\begin{document}
\Draw
   \Scale(1,0.5)
   \PieChartSpec(1,50,20)()
   \PieChart(15 & 10 & 30)
   \Move(0,-6)
   \DrawOvalArc(50,50)(180,360)
\EndDraw
\end{document}

\underline{latex sec161.Changeable} Chemical diagram in section 16.1 (for LaTeX)

\Define\ChemEdge(3){{
  \MoveToNodeDir(#2,#1)
  \MarkLoc(b) \MoveToNodeDir(#1,#2)
  \MarkLoc(a) \DSeg\RotateTo(a,b)
  \Rotate(-90)\Q=#3; \Q/2; \Q-0.5;
  \Q*2.2;  \MoveF(\Val\Q)
  \Do(1,#3){ {\CSeg\Line(a,b)}
                 \MoveF(-2.2) } }}
\Do(1,6){ \MoveF(30)
  \Node(\DoReg)(--C--) \MoveF(20) \Node(x)(--H--)
  \ChemEdge(\DoReg,x,1) \MoveTo(0,0) \Rotate(60) }
\ChemEdge(1,2,2) \ChemEdge(2,3,1) \ChemEdge(3,4,2)
\ChemEdge(4,5,1) \ChemEdge(5,6,2) \ChemEdge(6,1,1)

\underline{latex FA.Changeable} Finite automaton in last example of chapter 17 (for LaTeX)

\NewCIRCNode(\StateNode,106,)

\NewCIRCNode(\AStateNode,106,103)
\Define\StateAt(3){ \MoveTo(#2,#3)
   \StateNode(#1)(--$q_{#1}$--)}
\Define\AStateAt(3){
   \MoveTo(#2,#3)
   \AStateNode(#1)(--$q_{#1}$--)}
\DiagramSpec(\StateAt&
       \AStateAt&\TransEdge)
\ArrowHeads(1)     \Diagram
   ( 0,0,0    & 1,50,50  &
     3,90,-50 & 4,20,-50 )(2,75,0)
   (0,1,a,b & 1,2,a, & 0,2, ,b & 2,2,0,b & 0,4,a,  & 4,3,a,b )
\CurvedEdgeAt(3,1,0,1,1,0)(20,0.3,0,0.5) \EdgeLabel[+](--a--)

\underline{latex CM.Changeable} Commutative diagram in last example of section 18.3 (for LaTeX)

\GridDiagramSpec()(\MyEdge)
\Define\L(4){,+#1..+#2\,L\,#3\,#4}
\Define\D(4){,+#1..+#2\,D\,#3\,#4}
\Define\MyEdge(5){
   \IF \EqText(#3,D) \THEN\EdgeSpec(D)
   \ELSE  \EdgeSpec(L)  \FI
   \IF  \EqText(#1,#2)  \THEN
      \RotateTo(#4) \CycleEdge(#1)
      \EdgeLabel(--$#5$--)
   \ELSE\Edge(#1,#2)
      \IF \EqText(,#4) \THEN \EdgeLabel(--$#5$--)
      \ELSE \EdgeLabel[#4](--$#5$--) \FI
   \FI}
\ArrowHeads(1)
\GridDiagram(3,3)()()({
& T \D(0,1,,f) \D(2,0,,h) & N \L(2,0,,\sigma)//
S \L(0,0,180,t) \L(-1,1,,f) \L(1,1,+,g\circ f)//
& W \L(0,1,,f) & X \D(0,0,-45,r)//})

\underline{latex sec202.Changeable} Second example in section 20.2 (for LaTeX)

\NewNode(\MyNode,\MoveToRect){
   \Move(-24,-8) \DrawRect(48,16)
   \Move(16,0) {\Line(0,16)}
   \Move(16,0) \Line(0,16)  }
\Define\MyEdge(2){
   \CSeg\FindSign(#1,#2)
   \EdgeAt(#1,\sign 0.66,0,#2,0,1)}
\Define\FindSign(2){
   \IF  \GtDec(#1,0)  \THEN
        \Define\sign{}
   \ELSE \Define\sign{-} \FI}
\TextNode(2){ \Text(--$#2$--)
   \IF \EqInt(#1,0) \THEN {

     \Move(-16,0) \Text(--$\Lambda$--)
     \Move(32,0) \Text(--$\Lambda$--)
   }\FI }
\ArrowHeads(1)
\TreeSpace(D,10,20)
\TreeSpec(\MyNode)()(\MyEdge)
\Tree()(2,2,A// 0,0,B&2,2,C// 2,2,D&0,0,E// 0,0,F&0,0,G//)

\underline{latex LR.Changeable} Last example in section 20.2 (for LaTeX)

\GridDiagramSpec(\RNode)(\MyEdge)
\GridSpace(10,10) \ArrowSpec(F,5,3,3)
\ArrowHeads(1) \CycleEdgeSpec(-30,20)
\NewNode(\RNode,\MoveToOval){
  \GetNodeSize   \Va+4;  \Vb+4;
  \SetMinNodeSize   \Move(0,3)
  \DrawOval(\Val\Va,\Val\Vb) }
\Define\MyEdge(3){
  \IF \EqText(#1,#2) \THEN
    \RotateTo(45) \LabelPos(+0.5;,3)
    \CycleEdge(#1)
  \ELSE  \DSeg\Q(#1,#2)
    \IF \GtDec(\Q,180) \THEN
      \LabelPos(+0.4;,3) \ELSE \LabelPos(0.5;,3)
    \FI \Edge(#1,#2) \FI \EdgeLabel(--$#3$--) }
\baselineskip=9pt       \Define\NEXT(3){,#1..#2\,#3}
\def\/#1/#2.#3;{\hbox{$\scriptstyle#1\rightarrow #2\cdotp#3$}~~}
\GridDiagram(3,3)()()({
&  \/S'/ S. ; \/A / S.a; $I_1$ \NEXT(+0,+1,a)
&  \/A / Sa.;            $I_2$
// \/S'/ .S ; \/S / .AS; \/  / .b  ;
   \/A / .Sa; \/  / .a ; $I_0$ \NEXT(-1,+1,S) \NEXT(+0,+1,A)                                 \NEXT(+1,+1,b) \NEXT(+1,+0,a)
&  \/S /A.S ; \/S /.AS  ; \/  /.b   ;
   \/A /.Sa ; \/  /.a   ; $I_3$ \NEXT(+0,+1,S) \NEXT(+0,+0,A)
&   \/A/ S.a ;         $I_4$  \NEXT(+1,+0,a)
//
   \/A/a.;    $I_5$
&  \/S/ b. ;   $I_6$
&  \/A/ Sa. ;         $I_7$
//
})

\underline{latex TM.Changeable} Last example in chapter 20 (for LaTeX)

\NewCIRCNode(\StateNode,104,)
\NewCIRCNode(\AStateNode,104,102)
\Define\StateAt(3){ \MoveTo(#2,#3)
   \StateNode(#1)(--$q_{#1}$--)}
\Define\AStateAt(3){\MoveTo(#2,#3)
   \AStateNode(#1)(--$q_{#1}$--)}

\LabelSpec(4){\PictLabel{
{ \Move(1,-5) \Line(8,20) }
{ \Move(4,10) { \EntryExit(1,0,0,0)
  \Text(--\strut$#1$--) }  \Move(6,0)
  \EntryExit(-1,0,0,0) \Text(--\strut$#3$--) }
{ \EntryExit( 1,0,0,0) \Text(--\strut$#2$--) }
\Move(6,0)  \EntryExit(-1,0,0,0) \Text(--\strut$#4$--) }}
\DiagramSpec(\StateAt&\AStateAt&\TransEdge)  \ArrowHeads(1)
\Diagram( 0,0,0 & 1,80,0 )(2,80,-60)(
0,0,90,{a,B,+1,a{,+}1}                 &
0,1,   {a,B,+1,a{,+}1}, {b,B,0,B{,-}1} &
1,2,   {\$,B,0,B{, }0},                )

ProTeX and AlProTex

NOTE. The LaTEX files of this part put the \AlProTex.... command after the \begin.... command. This is not a nice style for files.

\underline{latex exr2222a.Changeable} Exercise 22.2.2 (for LaTeX)

\AlProTex{awk,<<<>>>,@,%
  list,title,[]}
\catcode`\@=0
An Awk program for deriving
the distribution of words.
@<dist@><<<
BEGIN {print"words\t lines"}
@<count the words@>
END { i=0
      while (i<=m)
      {if (c[i]>0)
       {print i "\t " c[i]}
       i=i+1 }}
>>>
NF is a built-in variable
that holds the number of
words in the current input
line.
@<count the words@><<<
{ c[NF]=c[NF]+1
  {if (NF>m) m=NF} }
>>> \OutputCode@<dist@>

\underline{latex exr2222b.Changeable} Exercise 22.2.2 (for LaTeX)

\AlProTex{awk,<<<>>>,@,%
  list,title,[]}
An Awk program for deriving
the distribution of words.
\<dist\><<<
BEGIN {print"words\t lines"}
@<count the words@>
END { i=0
      while (i<=m)
      {if (c[i]>0)
       {print i "\t " c[i]}
       i=i+1 }}
>>>

NF is a built-in variable
that holds the number of
words in the current input
line.
\<count the words\><<<
{ c[NF]=c[NF]+1
  {if (NF>m) m=NF} }
>>> \OutputCode\<dist\>

\underline{latex exr2223.Changeable} Exercise 22.2.3 (for LaTeX)

\AlProTex{p,<<<>>>,%
           title,list,[]}
\<prog\><<<
PROGRAM prog;>>>
\<body\><<<
{ comment } >>>
\<prog\><<<
BEGIN
  \<body\>
END.>>>
\OutputCode\<prog\>

\underline{latex exr2224a.Changeable} Exercise 22.2.4 (for LaTeX)

\AlProTex{p,list,%
   <<<>>>,title,[]}
\<\><<<
PROGRAM prog;>>>
\<body\><<<
{ comment } >>>
\<\><<<
BEGIN
  \<body\>
END.>>>
\OutputCode\<\>

\underline{latex exr2224b.Changeable} Exercise 22.2.4 (for LaTeX)

\AlProTex{p,%
    <<<>>>,title,[]}
\<\><<<
PROGRAM prog;>>>
\ShowCode-\<\>
\<body\><<<
{ comment } >>>
\ShowCode\<body\>
\<\><<<
BEGIN
  \<body\>
END.>>>
\ShowCode\<\>
\OutputCode\<\>

\underline{latex exr2224c.Changeable} Exercise 22.2.4 (for LaTeX)

\AlProTex{p,%
   <<<>>>,title,[]}
\<\><<<
PROGRAM prog;>>>
\<body\><<<
{ comment } >>>
\<\><<<
BEGIN
  \<body\>
END.>>>
\OutputCode\<\>

\underline{latex exr2225a.Changeable} Exercise 22.2.5 (for LaTeX)

\AlProTex{p,<<<>>>,%
          title,list,[]}
\<myprog\><<<
PROGRAM prog;>>>
\<body\><<<
{ comment } >>>
\<myprog\><<<
BEGIN
  \<body\>

END.>>>
\OutputCode\<myprog\>

\underline{latex exr2225b.Changeable} Exercise 22.2.5 (for LaTeX)

\AlProTex{p,title,list,[]}
\<myprog\>{PROGRAM prog;}
\<body\>{ \{ comment \} }
\<myprog\>{BEGIN
  \<body\>
END.}
\OutputCode\<myprog\>

\underline{latex exr2226a.Changeable} Exercise 22.2.6 (for LaTeX)

\AlProTex{p,<<<>>>,[],%
             title,list}
\<myprog\><<<
PROGRAM prog;
BEGIN
  \<body\>
END.>>>
\<body\><<<
{ comment: \\@|` } >>>

\underline{latex exr2226b.Changeable} Exercise 22.2.6 (for LaTeX)

\AlProTex{p,<<<>>>,[],%
             title,list,@}
\<myprog\><<<
PROGRAM prog;
BEGIN
  @<body@>
END.>>>
\<body\><<<
{ comment: \@@|` } >>>

\underline{latex exr2226c.Changeable} Exercise 22.2.6 (for LaTeX)

\AlProTex{p,<<<>>>,[],%

             title,list,|}
\<myprog\><<<
PROGRAM prog;
BEGIN
  |<body|>
END.>>>
\<body\><<<
{ comment: \@||` } >>>

\underline{latex exr2226d.Changeable} Exercise 22.2.6 (for LaTeX)

\AlProTex{p,<<<>>>,[],%
             title,list,`}
\<myprog\><<<
PROGRAM prog;
BEGIN
  `<body`>
END.>>>
\<body\><<<
{ comment: \@|`` } >>>

\underline{latex exr2227.Changeable} Exercise 22.2.7 (for LaTeX)

\AlProTex{p,<<<>>>,[],%
               title,list,|}
\<temp\><<<
\<prog\><<<
PROGRAM prog;
BEGIN
  ||<body||>
END.>|empty>>
\<body\><<<
{comment: \@||||} >|empty>>
\OutputCode\<prog\>  >>>
\OutputCode[prg]\<temp\>
\input temp.prg

\underline{latex exr2228a.Changeable} Exercise 22.2.8 (for LaTeX)

\Code\myprog{my
   program}{PROGRAM prog;}
\ShowCode-\myprog

\AppendCode\myprog{BEGIN
  \body
END.}
\Code\body{body}{}
\ShowCode-\myprog
\AppendCode\body{\{comment\}}
\ShowCode\body
Code of title \myprog{}
linearized within
the paragraph:
`\ShowCode-\myprog'.
\OutputCode\myprog

\underline{latex exr2228b.Changeable} Exercise 22.2.8 (for LaTeX)

\AlProTex{p,<<<>>>,list,@}
\Code\myprog{my program}<<<
PROGRAM prog;>>>
\Code\body{body}<<<
>>>
\AppendCode\myprog<<<
BEGIN
  @body
END.>>>
\AppendCode\body<<<
 {comment} >>>
Code of title \myprog{}
linearized within the
paragraph:
`\ShowCode-\myprog'.
\OutputCode\myprog

\underline{latex exr2229.Changeable} Exercise 22.2.9 (for LaTeX)

\AlProTex{p,<<<>>>,%
          title,list,[]}
\<myprog\><<<
PROGRAM prog;>>>
\<body\><<<

{ comment } >>>
\<myprog\><<<
BEGIN
  \<body\>
END.>>>
\OutputCode\<myprog\>

\underline{latex exr22210a.Changeable} Exercise 22.2.10 (for LaTeX)

\AlProTex{p,<<<>>>,title,%
         list,[],enumerate}
See lines \RefLine[*].
\<myprog\><<<
PROGRAM prog;\TagLine[*] >>>
\<body\><<<
{comment: see \RefLine[*]}>>>
\<myprog\><<<
BEGIN
  \<body\>   \TagLine[*] >>>
\ResetLineCount
\<myprog\><<<
END.>>> \OutputCode\<myprog\>

\underline{latex exr22210b.Changeable} Exercise 22.2.10 (for LaTeX)

\AlProTex{p,<<<>>>,title,%
                 list,[]}
\def\CodeRef[#1]{\Ref{#1}}
\def\CodeTag[#1][#2]{%
              \Tag{#1}{#2}}
See lines \Ref{*}.
\<myprog\><<<
PROGRAM prog;\CodeTag[*][1]>>>
\<body\><<<
{comment: see \CodeRef[*]}>>>
\<myprog\><<<
BEGIN
  \<body\>   \CodeTag[*][2]>>>

\ResetLineCount \Tag{*}{3}
\<myprog\><<<
END.>>> \OutputCode\<myprog\>

\underline{latex exr22211.Changeable} Exercise 22.2.11 (for LaTeX)

\AlProTex{p,<<<>>>,title,%
         list,[],ShowIndex}
\<myprog\><<<
PROGRAM prog;>>>
\<body\><<<
{ comment } >>>
\TextIndex{Tracing}
\SecIndex{***}
\<myprog\><<<
BEGIN
  \<body\>
END.>>>
\def\BeforeIndex{Index}
\ShowIndex

\underline{latex exr22212.Changeable} Exercise 22.2.12 (for LaTeX)

\AlProTex{p,<<<>>>,%
     title,[],ShowTop}
\<my prog\><<<
PROGRAM prog;
BEGIN  \<body\>
END.>>>
\<body\><<<
{ comment 1}
>>>
\<body\><<<
\<subbody\>
>>>
\<subbody\><<<
{ comment 2 } >>>
\ShowTop\<body\>
\def\ShowSep{(* ***** *)\par}
\ShowTop\<body\>

\underline{latex exr22213.Changeable} Exercise 22.2.13 (for LaTeX)

\AlProTex{p,<<<>>>,title,%
            list,[],NoShow}
First piece of text.
\<myprog\><<<
PROGRAM prog; >>>
\ShowOff Second piece of text.
\<body\><<<
{ comment } >>>
\Note  Third piece of
text. \EndNote\ShowOn
\<myprog\><<<
BEGIN
  \<body\>
END.>>>
Last piece of text.

\underline{latex exr22214.Changeable} Exercise 22.2.14 (for LaTeX)

\AlProTex{p,<<<>>>,title,%
   list,ShowIndex,ClearCode}
Chapter 1
\<progA\><<<
PROGRAM prog;
>>>
\<progA\><<<
BEGIN
  { comment 1 }
END.>>>    \ShowIndex
            \ClearCode
\vfil\break Chapter 2
\<progB\><<<
PROGRAM prog;
>>>
\<progB\><<<
BEGIN
  { comment 2 }
END.>>>  \ShowIndex

\underline{latex exr22215.Changeable} Exercise 22.2.15 (for LaTeX)

\hsize=2.5in \noindent{\bf
LITERATE PROGRAM IN  WEB}
\medskip  We compute a
two-dimensional view of a
three-dimensional point.
@* The mapping.  The view
   {\tt (x,y)} is  on the
   {\tt Z=0} plane.
@<map(x,y) := mapping of
   (X,Y,Z)@>=
   x := X * D DIV (D + Z);
@  The viewer is located
   at {\tt (0,0,-D)}.
@<CONST + VAR @>=
   CONST D=100;
   VAR  X {x}, Z: INTEGER;
@
@<map...@>=
   y := Y * D DIV (D + Z);
@
@<CON...@>=Y {y}: INTEGER;
@* Main program.
@p PROGRAM D2view;
     @<CON...@>
   BEGIN  read(X,Y,Z);
     @<map...@>
     write(x,y)
   END.

On a Unix system containing the xdvi application, the instructions of the following file can be executed with the compound command ‘chmod 777 execute.me ; execute.me’.

tex  sec62a.Changeable   # First example in section 6.2
xdvi  sec62a.dvi
latex  sec62b.Changeable   # First example in section 6.2
xdvi  sec62b.dvi
latex  sec62c.Changeable   # First example in section 6.2---second version
xdvi  sec62c.dvi
latex  exr672.Changeable   # Exercise 6.7.2
xdvi  exr672.dvi
tex  sec61.Changeable   # First example in section 6.1
xdvi  sec61.dvi
latex  sec61a.Changeable   # First example in section 6.1
xdvi  sec61a.dvi
latex  sec61b.Changeable   # First example in section 6.1---second version
xdvi  sec61b.dvi
latex  exr732.Changeable   # Exercise 7.3.2
xdvi  exr732.dvi
latex  exr741.Changeable   # Exercise 7.4.1
xdvi  exr741.dvi
latex  sec81.Changeable   # First example in section 8.1
xdvi  sec81.dvi
latex  usa.Changeable   # Example of USA map in section 9.3
xdvi  usa.dvi
latex  exr1011a.Changeable   # Exercise 10.1.1
xdvi  exr1011a.dvi
latex  exr1011b.Changeable   # Exercise 10.1.1
xdvi  exr1011b.dvi
latex  exr1011c.Changeable   # Exercise 10.1.1
xdvi  exr1011c.dvi
latex  exr1011d.Changeable   # Exercise 10.1.1
xdvi  exr1011d.dvi
latex  exr1011e.Changeable   # Exercise 10.1.1
xdvi  exr1011e.dvi
latex  exr1011f.Changeable   # Exercise 10.1.1
xdvi  exr1011f.dvi
latex  exr1021.Changeable   # Exercise 10.2.1
xdvi  exr1021.dvi
latex  exr1022.Changeable   # Exercise 10.2.2
xdvi  exr1022.dvi
latex  boxes.Changeable   # Last example in section 11.2
xdvi  boxes.dvi
latex  exr1121.Changeable   # Exercise 11.2.1
xdvi  exr1121.dvi
latex  exr1124.Changeable   # Exercise 11.2.4
xdvi  exr1124.dvi
latex  exr1131.Changeable   # Exercise 11.3.1
xdvi  exr1131.dvi
latex  exr1132.Changeable   # Exercise 11.3.2
xdvi  exr1132.dvi
tex  sec132.Changeable   # Example in section 13.2
xdvi  sec132.dvi
tex  sec132a.Changeable   # Example in section 13.2---second version

xdvi  sec132a.dvi
latex  sec132b.Changeable   # Example in section 13.2
xdvi  sec132b.dvi
latex  sec132c.Changeable   # Example in section 13.2---second version
xdvi  sec132c.dvi
latex  sec132d.Changeable   # Example in section 13.2---third version
xdvi  sec132d.dvi
latex  sec132e.Changeable   # Example in section 13.2---forth version
xdvi  sec132e.dvi
latex  sec161.Changeable   # Chemical diagram in section 16.1
xdvi  sec161.dvi
latex  FA.Changeable   # Finite automaton in last example of chapter 17
xdvi  FA.dvi
latex  CM.Changeable   # Commutative diagram in last example of section 18.3
xdvi  CM.dvi
latex  sec202.Changeable   # Second example in section 20.2
xdvi  sec202.dvi
latex  LR.Changeable   # Last example in section 20.2
xdvi  LR.dvi
latex  TM.Changeable   # Last example in chapter 20
xdvi  TM.dvi
latex  exr2222a.Changeable   # Exercise 22.2.2
xdvi  exr2222a.dvi
latex  exr2222b.Changeable   # Exercise 22.2.2
xdvi  exr2222b.dvi
latex  exr2223.Changeable   # Exercise 22.2.3
xdvi  exr2223.dvi
latex  exr2224a.Changeable   # Exercise 22.2.4
xdvi  exr2224a.dvi
latex  exr2224b.Changeable   # Exercise 22.2.4
xdvi  exr2224b.dvi
latex  exr2224c.Changeable   # Exercise 22.2.4
xdvi  exr2224c.dvi
latex  exr2225a.Changeable   # Exercise 22.2.5
xdvi  exr2225a.dvi
latex  exr2225b.Changeable   # Exercise 22.2.5
xdvi  exr2225b.dvi
latex  exr2226a.Changeable   # Exercise 22.2.6
xdvi  exr2226a.dvi
latex  exr2226b.Changeable   # Exercise 22.2.6
xdvi  exr2226b.dvi
latex  exr2226c.Changeable   # Exercise 22.2.6
xdvi  exr2226c.dvi
latex  exr2226d.Changeable   # Exercise 22.2.6
xdvi  exr2226d.dvi
latex  exr2227.Changeable   # Exercise 22.2.7
xdvi  exr2227.dvi
latex  exr2228a.Changeable   # Exercise 22.2.8
xdvi  exr2228a.dvi
latex  exr2228b.Changeable   # Exercise 22.2.8
xdvi  exr2228b.dvi
latex  exr2229.Changeable   # Exercise 22.2.9
xdvi  exr2229.dvi

latex  exr22210a.Changeable   # Exercise 22.2.10
xdvi  exr22210a.dvi
latex  exr22210b.Changeable   # Exercise 22.2.10
xdvi  exr22210b.dvi
latex  exr22211.Changeable   # Exercise 22.2.11
xdvi  exr22211.dvi
latex  exr22212.Changeable   # Exercise 22.2.12
xdvi  exr22212.dvi
latex  exr22213.Changeable   # Exercise 22.2.13
xdvi  exr22213.dvi
latex  exr22214.Changeable   # Exercise 22.2.14
xdvi  exr22214.dvi
latex  exr22215.Changeable   # Exercise 22.2.15
xdvi  exr22215.dvi